OpenStudy (anonymous):

How can this be factored a^2-2ab-8b^2

Directrix (directrix):

@marird15 Do you know how to factor a^2 - 2a - 8? I realize that is not the question asked but the factoring technique is similar.

Directrix (directrix):

Great. Would you post that here? Then, we can relate it to your question?

OpenStudy (anonymous):

(a-4)(a+2) ^^Post this up?

Directrix (directrix):

Correct. Look at your work: a^2 - 2a - 8 = (a-4)(a+2) Compare that to the trinomial you want to factor: a^2-2ab-8b^2. How do they differ. Compare term by term.

OpenStudy (anonymous):

the second one has two variables A and B

Directrix (directrix):

Yes. Think about the problem as: (a)^2 - 2 (a)(b) - 8 (b)^2 -4 and 2 will still be part of the factorization but will have to have the variable b with them. Any ideas on how to do that? Play with it a bit.

Directrix (directrix):

Hint: (a - ? ) * (a + ?) = a^2-2ab-8b^2

Directrix (directrix):

You know you need the -4 and the 2. You also need to account for the variable b.

Directrix (directrix):

What does -4b times 2b equal?

OpenStudy (anonymous):

- 8b wouldn't it be (a-4b)(a+2b) ?

Directrix (directrix):

-4b times 2b equals - 8b^2. And, yes, (a-4b)(a+2b) = a^2-2ab-8b^2 Very good.

OpenStudy (anonymous):

Yaay! okay thank u ! i understand now

Directrix (directrix):

Great. Thanks for hanging in there with me as we went through the process. It was fun.

Directrix (directrix):

Question? @marird15 Or, another problem to work?

OpenStudy (anonymous):

oh no thanks that was it! :) @Directrix